# ThinkAct: Vision-Language-Action Reasoning via Reinforced Visual Latent Planning

### v1

@@ 3.1 Problem Formulation
We first define the setting and notations for vision-language-action (VLA) reasoning tasks. At each timestep $t$, the model receives a visual observation $o_{t}$ and a textual instruction $l$, with the goal of predicting an action $a_{t}$, which can be a textual command or a 7-DOF control vector $\left[\Delta_{x},\Delta_{\theta},\Delta_{\text{Grip}}\right]$ depending on the embodiment. To tackle this problem, we propose ThinkAct, a unified framework that aims to leverage an MLLM $\mathcal{F}_{\theta}$ to reason the high-level plans while connecting with an action model $\pi_{\phi}$ to infer executable actions. The MLLM $\mathcal{F}_{\theta}$ produces a visual plan latent $c_{t}$ based on $(o_{t},l)$, capturing the high-level intent and planning context (Sec. 3.2). This reasoned plan $c_{t}$ then guides the downstream action module $\pi_{\phi}$ to sequentially predict $N$ executable actions $[a_{t}]_{t}^{t+N}$ tailored to the target environment (Sec. 3.3). By connecting abstract planning with low-level control, our ThinkAct enables long-horizon reasoning and improves action adaptation in dynamic embodied tasks.
@@ 3.2 Reinforced Visual Latent Planning for Embodied Reasoning
$r_{\text{traj}}=\max\left(0,1-d(\tau,\hat{\tau})\right).$ (2)
Here, $d(\tau,\hat{\tau})$ denotes a metric measuring the distance between two trajectories, i.e., dynamic time warping (DTW) distance Senin (2008) in this work.
@@
$\ \text{where}\quad A_{i}=\frac{r_{i}-\text{mean}(\{r_{1},\ldots,r_{M}\})}{\text{std}(\{r_{1},\ldots,r_{M}\})}.$
@@ 4.2 Quantitative Evaluation
Dataset Split Octo-Base RT1-X OpenVLA DiT-Policy TraceVLA CoT-VLA Magma ThinkAct (Ours)
Simpler-Google (Visual Matching) Open/Close Drawer 1.0 22.5 49.5 44.9 57.0 56.0 50.0
@@
Simpler-Google (Variant Aggregation) Open/Close Drawer 22.0 56.0 22.5 35.5 31.0 53.4 47.6
@@
Simpler-Bridge (Visual Matching) Put Carrot on Plate 8.3 4.2 4.2 29.4 31.0 37.5
@@
Dataset Split / Metric GPT-4V LLaVA-Video InternVL2.5 InternVL3 NVILA Qwen2.5-VL Qwen2.5-VL* Magma ThinkAct (Ours)
EgoPlan- Bench2 Daily life 36.7 38.0 36.2 38.5 35.8 31.4 47.9 32.1 50.1
@@ B.4 Results of Smaller Model Size
Dataset Split / Metric InternVL2.5-2B InternVL3-2B NVILA-2B Qwen2.5-VL-3B Qwen2.5-VL-3B* ThinkVLA-3B (Ours)
EgoPlan- Bench2 Daily life 30.9 36.9 34.6 29.0 44.9 46.6
@@ B.6 Ablation Study
We ablate the frequency of reasoning updates by varying the number of actions per reasoning step $N$ on LIBERO. Setting $N$ to 25, 50, 75, and 100 results in average success rates of 84.0%, 84.6%, 84.4%, and 83.7%, respectively. These results suggest that overly sparse reasoning (e.g., $N{=}100$) might cause the model to be unable to detect the failure and perform self-correction in time, leading to degraded performance. On the other hand, too frequent updates (e.g., $N{=}25$) would induce additional inference cost without yielding substantial performance gains. As a result, we set the number of actions per reasoning $N$ as 75 on LIBERO.

### v2

@@ 3.1 Problem Formulation
We first define the setting and notations for vision-language-action (VLA) reasoning tasks. At each timestep $t$, the model receives a visual observation $o_{t}$ and a textual instruction $l$, with the goal of predicting an action $a_{t}$, which can be a textual command or a 7-DOF control vector $\left[\Delta_{x},\Delta_{\theta},\Delta_{\text{Grip}}\right]$ depending on the embodiment. To tackle this problem, we propose ThinkAct, a unified framework that aims to leverage an MLLM $\mathcal{F}_{\theta}$ to reason the high-level plans while connecting with an action model $\pi_{\phi}$ to infer executable actions. The MLLM $\mathcal{F}_{\theta}$ produces a visual plan latent $c_{t}$ based on $(o_{t},l)$, capturing the high-level intent and planning context (Sec. 3.2). This reasoned plan $c_{t}$ then guides the downstream action module $\pi_{\phi}$ to sequentially predict $N$ executable actions $a_{t}]_{t}^{t+N}$ tailored to the target environment (Sec. 3.3). By connecting abstract planning with low-level control, our ThinkAct enables long-horizon reasoning and improves action adaptation in dynamic embodied tasks.
@@ 3.2 Reinforced Visual Latent Planning for Embodied Reasoning
[…]
@@
[…]
@@ 4.2 Quantitative Evaluation
Dataset Split
Octo-Base
RT1-X
OpenVLA
DiT-Policy
TraceVLA
CoT-VLA
Magma
ThinkAct (Ours)
Simpler-Google
(Visual Matching)
Open/Close Drawer 1.0 22.5 49.5 44.9 57.0 56.0 50.0
@@
Simpler-Google
(Variant Aggregation)
Open/Close Drawer 22.0 56.0 22.5 35.5 31.0 53.4 47.6
@@
Simpler-Bridge
(Visual Matching)
Put Carrot on Plate 8.3 4.2 4.2 29.4 31.0 37.5
@@
Dataset Split / Metric
GPT-4V
LLaVA-Video
InternVL2.5
InternVL3
NVILA
Qwen2.5-VL
Qwen2.5-VL*
Magma
ThinkAct
EgoPlan-
Bench2
Daily life 36.7 38.0 36.2 38.5 35.8 31.4 47.9 32.1 50.1
@@ B.4 Results of Smaller Model Size
Dataset Split / Metric
InternVL2.5-2B
InternVL3-2B
NVILA-2B
Qwen2.5-VL-3B
Qwen2.5-VL-3B*
ThinkAct-3B
EgoPlan-
Bench2
Daily life 30.9 36.9 34.6 29.0 44.9 46.6
@@ B.6 Ablation Study
We ablate the frequency of reasoning updates by varying the number of actions per reasoning step $N$ on LIBERO. Setting $N$ to 25, 50, 75, and 100 results in average success rates of 84.0%, 84.6%, 84.4%, and 83.7%, respectively. These results suggest that overly sparse reasoning (e.g., N=100N{=}100) might cause the model to be unable to detect the failure and perform self-correction in time, leading to degraded performance. On the other hand, too frequent updates (e.g., N=25N{=}25) would induce additional inference cost without yielding substantial performance gains. As a result, we set the number of actions per reasoning $N$ as 75 on LIBERO.